OpenStudy (anonymous):

Using the square root property 6x squared =25. The answer is given, - + 5 square root 6/6. but I'm not sure how to work it.

OpenStudy (anonymous):

I know I square both sides, but I don't come up with the right answer.

OpenStudy (anonymous):

can u show me your steps so I can help you fix them

OpenStudy (anonymous):

\[6x ^2=25\] I divided both sides by 6 then squared both sides, I got \[x = 5/6\]

OpenStudy (anonymous):

let me see what I get real quick

OpenStudy (anonymous):

\[+-5\sqrt6\div6\] is the answer given. thank you

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hold up im still solving it

OpenStudy (anonymous):

holding :)

OpenStudy (anonymous):

My example: (follow these steps) 4x^2 - 20x + 25 = 7 (factor) (2x - 5)^2 = 7 (square root both sides) 2x - 5 = ±√7 (add 5 to both sides) 2x = 5 ± √7 (divide both sides by 2) x = (5 ± √7)/2 <==== then you get your answer. :D

OpenStudy (anonymous):

or you can try this one to help you x^2 + 6x + 9 = 25 (factor) (x + 3)^2 = 25 (square root both sides) x + 3 = ±5 (subtract 3 from both sides) x = -3 ± 5 x = -3 - 5 or x = -3 + 5 x = -8 or x = 2

OpenStudy (anonymous):

ok..thank you for your time!

OpenStudy (anonymous):

anytime you need help you can come to me